# Anticipating Gaming to Incentivize Improvement: Guiding Agents in (Fair) Strategic Classification

### v1

@@ 3 Problem Setting and Preliminaries
The firm.
A firm makes decisions on agents based on their observable features $x$ and their group memberships $s$. The decision is binary, denoted $d\in\{0,1\}$, where $d=1$ represents acceptance and $d=0$ represents rejection. This decision is determined by a group-dependent binary classifier $\pi_{s}(x)=\mathbb{P}(D=1|X=x,S=s)$. We assume that this is a threshold policy, such that an agent is accepted if and only if ${x}\geq\theta_{s}$. (Zhang et al. [2020b] show that such threshold policies are optimal under mild assumptions.)
Agents’ strategic actions.
@@
Taking these actions may impact the agents’ features and/or true labels. We use $X$ and $Y$ to denote the pre-strategic feature and true label (before an action $w$ is taken), and $\hat{X}$ and $\hat{Y}$ to denote the post-strategic ones. In particular, both qualified and unqualified agents ($y=1$ and $y=0$) can opt to manipulate ($w=M$) by changing their feature $x$ to some new feature $\hat{x}$, without changing their true qualification state ($\hat{y}=y$); e.g., this could represent cheating on an exam, which increases the score without affecting true learning. Alternatively, both types of agents can choose to invest in improving themselves ($w=I$), which leads to a change in both their feature from $x$ to some $\hat{x}$, as well as their true label to $\hat{y}=1$; e.g., this could represent studying for an exam, which improves both the score and true learning. Lastly, agents who opt to do nothing ($w=N$) maintain their feature and label ($\hat{x}=x$ and $\hat{y}=y$). The firm will only observe the agents’ post-strategic features $\hat{x}$ when making its decision. We let $\hat{\alpha}_{s}$ and $\hat{G}_{s}^{y}(x)$ denote the post-strategic population statistics after these changes have happened as a result of agents’ selected strategic actions. (We will formally characterize these updated statistics in Section 4.2).
In addition to differing in their impacts on changes in features and/or true labels, the actions differ in two aspects: their cost and their efficacy. First, each action requires exerting effort and comes at a certain group-dependent (constant) cost ${C_{w,s}}\in[0,1)$, $w\in\{M,I,N\},s\in\{a,b\}$. Such difference in costs across actions and groups is a common assumption in prior works on strategic classification (e.g., Zhang et al. [2022], Liu et al. [2020], Jin et al. [2022]). The element that is new to our model is that we assume all actions lead to a (weak) increase in the feature (i.e., we assume that ${\hat{x}\geq x}$), but that this increase in feature is stochastic and different across actions and groups. Formally, we assume the probability that the feature $x$ of a qualified/unqualified individual from a group $s$ increases by $b_{w}$ after opting for action $w$ is distributed according to a boost distribution with pdf ${\mathbb{\tau}^{y}_{w,s}(b)}:=\mathbb{P}(\hat{X}=x+b|X=x,Y=y,W=w,S=s)$. These boost distributions determine the efficacy of the action. As an example, they could capture that both cheating and studying can increase an agent’s exam score, but that the efficacy of these actions in improving scores is different. Let $\{\underline{b}^{y}_{w,s},\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu^{y}_{w,s}\}$ denote the minimum and maximum boost under action $w$, and $\mathbb{T}^{y}_{w,s}(b)$ denote the CDF of the boost function.
@@
In words, an action first-order stochastically dominating another means that it “gets the agent further”, in that it has a higher probability of increasing the agent’s feature from $x$ to a feature greater or equal to $\hat{x}=x+b$. Assumption 2 gives rise to the conflict between the two actions: manipulation is cheaper, while improvement is more effective in advancing the agent. (We note that there are three scenarios beyond Assumption 2: one can be recovered by interchanging the indexes $M$ and $I$ in our results. The remaining two cases are less interesting, as one of the actions would be more beneficial in both cost and efficacy.)
Agents’ utility.
@@
Formally, an agent chooses to incur the cost $C_{w,s}$ of action $w$ if and only if it (sufficiently) increases the probability that the agent is accepted by the firm. For an agent from group $s$ with feature $x$ and true label $y$, the benefit from strategic action $w$ is the increase in the probability of being accepted, given by
@@
The fairness constraint here, donated by $f$, will be considered to be one of the two more commonly studied notions of fairness Mehrabi et al. [2021]:
@@ 4 Agents’ Strategic Behavior
We begin by characterizing agents’ best responses to a firm’s given classifier. Consider an agent from group $s$ with feature $x$ and true label $y$, facing decision threshold $\theta_{s}$. We will show that for any given $y$ and $\theta_{s}$, the feature space $x$ can be partitioned into disjoint regions determining the agents’ best-response, with the boundaries of these regions determined by the points where agents become indifferent between pairs of actions.
@@
Type 1 $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$ $[0,\mathbf{o}^{y}_{I,s}):N$, $[\mathbf{o}^{y}_{I,s},\mathbf{r}^{y}_{s}):I$, $[\mathbf{r}^{y}_{s},\theta_{s}):M$, $[\theta_{s},\infty):N$
Type 2 $\mathbf{o}^{y}_{M,s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{f}^{y}_{s}$ $[0,\mathbf{o}^{y}_{M,s}):N$, $[\mathbf{o}^{y}_{M,s},\mathbf{f}^{y}_{s}):M$, $[\mathbf{f}^{y}_{s},\mathbf{r}^{y}_{s}):I$, $[\mathbf{r}^{y}_{s},\theta_{s}):M$, $[\theta_{s},\infty):N$
Type 3 $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{o}^{y}_{I,s}$ $[0,\mathbf{o}^{y}_{M,s}):N$, $[\mathbf{o}^{y}_{M,s},\theta_{s}):M$, $[\theta_{s},\infty):N$
@@
  • $[0,\mathbf{o}^{y}_{I,s}):N$, $[\mathbf{o}^{y}_{I,s},\mathbf{f}^{y}_{s}):I$, $[\mathbf{f}^{y}_{s},\theta_{s}):M$, $[\theta_{s},\infty):N$.
A detailed proof is given in Appendix A.1.1. These three types of best-responses are illustrated in Figure 1. In particular, note that in all types of equilibrium, agents who are close to the decision threshold opt to be risk takers, choosing uncertain but cheap manipulation over certain but costly improvement. Interestingly, we posit that this may present a parallel with gaming behavior observed in education settings: students committing academic dishonesty typically have higher GPAs [OEDB, n.d].
@@
We also note that the type of equilibrium is determined solely by the relative cost and efficacy of the manipulation and improvement actions; the choice of the classifier $\theta_{s}$ changes the indifference points where agents opt for each action, but not the type of equilibrium. We provide additional intuition for this below.
4.1 Illustration: uniform boost distributions
[…]
@@ 4.2 Post-strategic population statistics
We next characterize the post-strategic population statistics. In particular, as noted earlier, the strategic actions $w\in\{M,I\}$ change the agent’s true qualification and/or feature. Denote the post-strategic qualification rates by $\hat{\alpha}_{s}=\mathbb{P}(\hat{Y}=1|S=s)$, and the post-strategic feature distribution by ${\hat{G}^{y}_{s}(x)}:=\mathbb{P}(\hat{X}={x}|\hat{Y}={y},S=s)$. Let $\mathbb{I}^{y}_{s,(i)}:=\{x|Y=y,S=s,w=I\}$ be the set of agents from group $s$ with label $y$ who opt for improvement under a given threshold policy $\theta_{s}$, in a type (i) best response (as identified in Proposition 2). Define $\mathbb{M}^{y}_{s,(i)}$ and $\mathbb{N}^{y}_{s,(i)}$ similarly for the set of agents who opt for manipulation and for doing nothing under equilibrium type (i).
The following lemma presents the expressions for the post-strategic population statistics in terms of the pre-strategic population statistics and the best-response regions.
@@
$\hat{G}^{0}_{s}(x)$ $=\frac{1-\alpha_{s}}{1-\hat{\alpha}_{s}}\Big{(}\mathbbm{1}(x\in\mathbb{N}^{0}_{s,(i)})G^{0}_{s}(x)+(G^{0}_{s}*\tau^{0}_{M,s})(x)\Big{)},$ (5)
$\hat{G}^{1}_{s}(x)$ $=\frac{\alpha_{s}}{\hat{\alpha}_{s}}\Big{(}\mathbbm{1}(x\in\mathbb{N}^{1}_{s,(i)})G^{1}_{s}(x)+(G^{1}_{s}*\tau^{1}_{M,s})(x)+(G^{1}_{s}*\tau^{1}_{I,s})(x)\Big{)}+\frac{1-\alpha_{s}}{\hat{\alpha}_{s}}(G^{0}_{s}*\tau^{0}_{I,s})(x),$ (6)
@@ 5 The Firm’s Optimal Policy
This is similar to results obtained in [Zhang et al., 2022, Liao and Naghizadeh, 2022]. The proof is provided in Appendix A.2.1 for completeness.
@@
$\mathbb{E}[\hat{U}({\theta}_{a},{\theta}_{b})]$ $:=\mathbb{E}[U({\theta}_{a},{\theta}_{b})]$
$+\sum_{s\in\{a,b\}}n_{s}(u_{+}\alpha_{s}\bm{\Phi}_{s,(i)}^{1}({\theta_{s}})+u_{+}(1-\alpha_{s})\bm{\Phi}_{s,(i)}^{0}({\theta_{s}})+u_{+}\alpha_{s}\bm{\Psi}_{s,(i)}^{1}({\theta_{s}})-u_{-}(1-\alpha_{s})\bm{\Psi}_{s,(i)}^{0}({\theta_{s}})),$ (7)
@@
The unfair strategic firm’s optimal decision thresholds $\hat{\theta}_{s}^{U}$ for best-response type (i) satisfy $\frac{G_{s}^{1}(\hat{\theta}^{U}_{s})-\bm{\Phi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta}^{U}_{s})-\bm{\Psi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta}^{U}_{s})-\frac{(1-\alpha_{s})}{\alpha_{s}}\bm{\Phi_{s,(i)}^{{}^{\prime}0}}(\hat{\theta}^{U}_{s})}{G_{s}^{0}(\hat{\theta}^{U}_{s})-\bm{\Psi_{s,(i)}^{{}^{\prime}0}}(\hat{\theta}^{U}_{s})}=\frac{u_{-}(1-\alpha_{s})}{u_{+}\alpha_{s}}$, where $\bm{\Phi_{s,(i)}^{{}^{\prime}y}}$ (resp. $\bm{\Psi_{s,(i)}^{{}^{\prime}y}}$) is the first order derivative of $\bm{\Phi_{s,(i)}^{y}}$ (resp. $\bm{\Psi_{s,(i)}^{y}}$) with respect to ${\theta_{s}}$, and with the specific forms under each best-response type given in Table 3.
@@
The fair non-strategic firm’s optimal decision thresholds $\theta_{s}^{\mathcal{C}}$ satisfy:
$\sum_{s}n_{s}\frac{u_{+}\alpha_{s}G^{1}_{s}(\theta_{s}^{\mathcal{C}})-u_{-}(1-\alpha_{s})G^{0}_{s}(\theta_{s}^{\mathcal{C}})}{\frac{\partial{\mathcal{C}^{f}_{s}(\theta_{s}^{\mathcal{C}})}}{\partial{\theta_{s}^{\mathcal{C}}}}}=0$.
@@
The fair strategic firm’s optimal decision thresholds $\hat{\theta}_{s}^{\mathcal{C}}$ satisfy:
$\sum_{s}n_{s}$ $\frac{u_{+}\alpha_{s}(\bm{\Phi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta_{s}^{\mathcal{C}}})+\bm{\Psi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta_{s}^{\mathcal{C}}})-G_{s}^{1}(\hat{\theta_{s}^{\mathcal{C}}}))+(1-\alpha_{s})(u_{+}\bm{\Phi_{s,(i)}^{{}^{\prime}0}}(\hat{\theta_{s}^{\mathcal{C}}})-u_{-}(\bm{\Psi_{s,(i)}^{{}^{\prime}0}(\hat{\theta_{s}^{\mathcal{C}}})}-G_{s}^{0}(\hat{\theta_{s}^{\mathcal{C}}})))}{\frac{\partial{\mathcal{C}^{f}_{s}(\hat{\theta}_{s}^{\mathcal{C}})}}{\partial{\theta_{s}^{\mathcal{C}}}}}=0$.
The proofs follow from applying similar techniques to those in [Zhang et al., 2022], combined with the results of Lemma 5; they are included in Appendix A.2.3 for completeness.
Lemmas 4-7, together with Proposition 2, fully characterize the possible (Stackelberg) equilibria of the game between agents and the firm, depending on the relative cost-efficacy of strategic actions, whether the firm is cognizant of agents’ strategic behavior, and whether the firm imposes any fairness constraints. We next use these characterizations to assess the impacts of anticipating agents’ strategic behavior and imposing fairness constraints on the optimal decision thresholds and the firm’s utility.
@@ 6 Effects of Anticipating Strategic Behavior
We first proceed with analyzing the impacts of anticipating agents’ strategic behavior on the optimal policies and the firm’s utility by comparing the non-strategic policy $\theta^{U}_{s}$ (from Lemma 4) with the strategic policy $\hat{\theta}_{s}$ (from Lemma 5) in the absence of any fairness interventions. For concreteness, we do this analysis for cases when all agents opt for the same type of best-response, and specifically, for Types 1 and 3 of agents’ best responses, which can be interpreted as relatively low and high improvement costs, respectively, as illustrated in Section 4.1. Accordingly, we will also refer to Type 1 and Type 3 equilibria as “affordable improvement” and “costly improvement”, respectively. Notably, Type 3/costly improvement equilibria only include manipulation decisions, whereas Type 1/affordable improvement equilibria include both manipulation and improvement decisions at equilibrium. Contrasting these two types of equilibria allows us to compare our findings with prior work, e.g. [Zhang et al., 2022], which have studied the (fairness) implications of anticipating gaming given only manipulation decisions. With our model, we can investigate to what extent anticipating gaming can prevent manipulation (as shown in prior work) but also potentially encourage improvement decisions (as we show here), and how the extent of each differs across qualification states and across groups.
@@
Assumption 4 states that the cost and effectiveness of the manipulation action is independent of the qualification state of agents. For instance, using a GPT tool to solve basic homework problems has the same cost and effectiveness regardless of a student’s preparation level/qualification state. In contrast, the second part of the assumption implies that improvement action is more effective for qualified agents. Continuing on the same example, when a qualified student with a stronger background spends time preparing well for an exam, their effort can boost their exam grades more than a less prepared student spending the same time.
Given Assumptions 3 and 4, we are now ready to establish relative orderings between $\theta_{s}$ and $\hat{\theta}_{s}$, and their impact on the different components of the firm’s utility. Recall, from the expression for the firm’s utility (5), that $\bm{\Phi}_{s,(i)}^{y}({\theta_{s}})$ and $\bm{\Psi}_{s,(i)}^{y}({\theta_{s}})$ capture the change in the firm’s utility due to label $y$ agents from group $s$ who opt for improvement and manipulation, respectively, in equilibrium Type (i), under a threshold $\theta_{s}$.
@@
    The strategic firm drives more improvement by both qualified and unqualified agents: $\bm{\Phi}^{y}_{s,(1)}(\theta^{U}_{s})<\bm{\Phi}^{y}_{s,(1)}(\hat{\theta}^{U}_{s}),~{}y\in\{0,1\}$.
The proofs are provided in Appendix A.3. We will next provide intuitive interpretations for the findings of this proposition. We discuss the impacts of relaxing the assumptions of Proposition 8 in Appendix A.4, as well as in the numerical experiments in Section 7.
6.1 Intuitive interpretation and key takeaways of Proposition 8
To highlight the intuition from Proposition 8, we will leverage the illustrations in Figure 2. In this figure, the horizontal, colored bars show the different regions of agents’ strategic best-responses under the costly improvement (Type 3) and affordable improvement (Type 1) equilibria (as shown in Proposition 2), for the non-strategic $\theta^{U}_{s}$ and strategic $\hat{\theta}^{U}_{s}$ policies. The curves above each plot show the pre-strategic feature distribution of agents with qualification state $Y$. The lightly shaded green/red ovals are used to further highlight the agents that opt for improvement/manipulation in each case.
(a) Costly improvement (Type 3) equilibrium
(b) Affordable improvement (Type 1) equilibrium
The costly improvement (Type 3) equilibrium.
The affordable improvement (Type 1) equilibrium.
Together, these observations highlights how anticipating strategic responses will not only reduce gaming of the algorithm (as also shown in prior work), but also lead to increased improvement incentives among agents (a new effect identified by our model).

### v2

@@ 3.1 The agents
3.2 The firm
A firm makes decisions on agents based on their observable features $x$ and their group memberships $s$. The decision is binary, denoted $d\in\{0,1\}$, where $d=1$ represents acceptance and $d=0$ represents rejection. This decision is determined by a group-dependent binary classifier $\pi_{s}(x)=\mathbb{P}(D=1|X=x,S=s)$. We assume that this is a threshold policy, such that an agent is accepted if and only if ${x}\geq\theta_{s}$. Zhang et al. [2020b] show that such threshold policies are optimal under mild assumptions.
3.3 Agents’ strategic actions
@@
Taking these actions may impact the agents’ features and/or true labels. We use $X$ and $Y$ to denote the pre-strategic feature and true label (before an action $w$ is taken), and $\hat{X}$ and $\hat{Y}$ to denote the post-strategic ones. In particular, both qualified and unqualified agents ($y=1$ and $y=0$) can opt to manipulate ($w=M$) by changing their feature $x$ to some new feature $\hat{x}$, without changing their true qualification state ($\hat{y}=y$); e.g., this could represent cheating on an exam, which increases the score without affecting true learning. Alternatively, both types of agents can choose to invest in improving themselves ($w=I$), which leads to a change in both their feature from $x$ to some $\hat{x}$, as well as their true label to $\hat{y}=1$; e.g., this could represent studying for an exam, which improves both the score and true learning. Lastly, agents who opt to do nothing ($w=N$) maintain their feature and label ($\hat{x}=x$ and $\hat{y}=y$). The firm will only observe the agents’ post-strategic features $\hat{x}$ when making its decision. We let $\hat{\alpha}_{s}$ and $\hat{G}_{s}^{y}(x)$ denote the post-strategic population statistics after these changes have happened as a result of agents’ selected strategic actions. (We analytically characterize these updated statistics as a function of their pre-strategic counterparts in Section 4.1).
In addition to differing in their impacts on changes in features and/or true labels, the actions differ in two aspects: their cost and their efficacy. First, each action requires exerting effort and comes at a certain group-dependent (constant) cost ${C_{w,s}}\in[0,1)$, $w\in\{M,I,N\},s\in\{a,b\}$. Such difference in costs across actions and groups is a common assumption in prior works on strategic classification (e.g., Zhang et al. [2022], Liu et al. [2020], Jin et al. [2022]). The element that is new to our model is that we assume all actions lead to a (weak) increase in the feature (i.e., we assume that $\hat{x}\geq x$}), but that this increase in feature is stochastic and different across actions and groups. Formally, we assume the probability that the feature $x$ of a qualified/unqualified individual from a group $s$ increases by $b_{w}$ after opting for action $w$ is distributed according to a boost distribution with pdf ${\mathbb{\tau}^{y}_{w,s}(b)}:=\mathbb{P}(\hat{X}=x+b|X=x,Y=y,W=w,S=s)$. These boost distributions determine the efficacy of the action. As an example, they could capture that both cheating and studying can increase an agent’s exam score, but that the studying is more effective for increasing scores. Let $\{\underline{b}^{y}_{w,s},\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu^{y}_{w,s}\}$ denote the minimum and maximum boost under action $w$, and $\mathbb{T}^{y}_{w,s}(b)$ denote the CDF of the boost function.
@@
In words, an action first-order stochastically dominating another means that it “gets the agent further”, in that it has a higher probability of increasing the agent’s feature from $x$ to a feature greater or equal to $\hat{x}=x+b$. Assumption 2 gives rise to the conflict between the two actions: manipulation is cheaper, while improvement is more effective in advancing the agent. (We note that there are three scenarios beyond Assumption 2: one can be recovered by interchanging the indexes $M$ and $I$ in our results. The remaining two cases are less interesting, as one of the actions would trivially dominate the other.)
3.4 Agents’ utility
@@
Formally, an agent chooses to incur the cost $C_{w,s}$ of action $w$ if and only if it (sufficiently) increases the probability of being accepted by the firm. For an agent from group $s$ with feature $x$ and true label $y$, the benefit from strategic action $w$ is the increase in the probability of being accepted, given by
@@ 3.5 The firm’s utility
The fairness constraint here, donated by $f$, will be considered to be one of the two more commonly studied notions of fairness [Mehrabi et al., 2021]:
@@ 4 Agents’ Strategic Behavior
We begin by characterizing agents’ best responses to a firm’s given classifier. We show that for any given qualification state $y$, group $s$, and corresponding threshold $\theta_{s}$, the feature space $x$ can be partitioned into disjoint regions determining the agents’ best-response, with the boundaries of these regions determined by the points where agents become indifferent between pairs of actions.
@@
Type 1 $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$ * $0,\mathbf{o}^{y}_{I,s}):N$, $\mathbf{o}^{y}_{I,s},\mathbf{r}^{y}_{s}):I$, $\mathbf{r}^{y}_{s},\theta_{s}):M$, $\theta_{s},\infty):N$
Type 2 $\mathbf{o}^{y}_{M,s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{f}^{y}_{s}$ $0,\mathbf{o}^{y}_{M,s}):N$, $\mathbf{o}^{y}_{M,s},\mathbf{f}^{y}_{s}):M$, $\mathbf{f}^{y}_{s},\mathbf{r}^{y}_{s}):I$, $\mathbf{r}^{y}_{s},\theta_{s}):M$, $\theta_{s},\infty):N$
Type 3 $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{o}^{y}_{I,s}$ $0,\mathbf{o}^{y}_{M,s}):N$, $\mathbf{o}^{y}_{M,s},\theta_{s}):M$, $\theta_{s},\infty):N$
@@
  • $0,\mathbf{o}^{y}_{I,s}):N$, $\mathbf{o}^{y}_{I,s},\mathbf{f}^{y}_{s}):I$, $\mathbf{f}^{y}_{s},\theta_{s}):M$, $\theta_{s},\infty):N$.
A detailed proof is given in Appendix A.1.1. These three types of best-responses are illustrated in Figure 1. In particular, note that in all types of equilibrium, agents who are close to the decision threshold opt to be risk takers, choosing uncertain but cheap manipulation over certain but costly improvement. Interestingly, we posit that this may present a parallel with gaming behavior observed in education settings: students committing academic dishonesty typically have higher GPAs [Open Education Database, n.d].
@@
We also note that the type of equilibrium is determined solely by the relative cost and efficacy of the manipulation and improvement actions; the choice of the classifier $\theta_{s}$ changes the indifference points where agents opt for each action, but not the type of equilibrium. Specifically, when the improvement cost is low, the best response is of Type 1, as agents find it beneficial to opt for improvement (higher efficacy, and relatively low cost) before manipulation becomes beneficial (which is near the threshold where risk of negative outcome is low enough). Moreover, when the improvement cost is Moderate, the best response is of Type 2, as now improvement is too expensive for agents who are far from the threshold, while still low enough that agents opt for improvement over manipulation when both actions are uncertain. Lastly, when the improvement cost is high, the best response is of Type 3, as now all agents choose the manipulation (once beneficial) over improvement, due to the considerably high improvement cost. We illustrate these intuitive interpretations more formally using uniform boost distributions in Appendix B.
@@ 4.1 Post-strategic population statistics
We next characterize the post-strategic population statistics. In particular, as noted earlier, the strategic actions $w\in\{M,I\}$ change the agent’s true qualification and/or feature. Denote the post-strategic qualification rates by $\hat{\alpha}_{s}=\mathbb{P}(\hat{Y}=1|S=s)$, and the post-strategic feature distribution by ${\hat{G}^{y}_{s}(x)}:=\mathbb{P}(\hat{X}={x}|\hat{Y}={y},S=s)$. Let $\mathbb{I}^{y}_{s,(i)}:=\{x|Y=y,S=s,w=I\}$ be the set of agents from group $s$ with label $y$ who opt for improvement under a given threshold policy $\theta_{s}$, in a type (i) best response (as identified in Proposition 2). Define $\mathbb{M}^{y}_{s,(i)}$ and $\mathbb{N}^{y}_{s,(i)}$ similarly for the set of agents who opt for manipulation and for doing nothing under equilibrium type (i). The following lemma presents the expressions for the post-strategic population statistics in terms of the pre-strategic population statistics and the best-response regions.
@@
$\hat{G}^{0}_{s}(x)$ $=\frac{1-\alpha_{s}}{1-\hat{\alpha}_{s}}\Big(\mathds{1}(x\in\mathbb{N}^{0}_{s,(i)})G^{0}_{s}(x)+(G^{0}_{s}*\tau^{0}_{M,s})(x)\Big),$ (5)
$\hat{G}^{1}_{s}(x)$ $=\frac{\alpha_{s}}{\hat{\alpha}_{s}}\Big(\mathds{1}(x\in\mathbb{N}^{1}_{s,(i)})G^{1}_{s}(x)+(G^{1}_{s}*\tau^{1}_{M,s})(x)+(G^{1}_{s}*\tau^{1}_{I,s})(x)\Big)+\frac{1-\alpha_{s}}{\hat{\alpha}_{s}}(G^{0}_{s}*\tau^{0}_{I,s})(x),$ (6)
@@ 5 The Firm’s Optimal Policy
This is similar to results obtained in prior work (e.g., [Zhang et al., 2022, Liao and Naghizadeh, 2022]). The proof is provided in Appendix A.2.1 for completeness.
@@
$\hskip-14.45377pt\mathbb{E}[\hat{U}({\theta}_{a},{\theta}_{b})]$ $:=\mathbb{E}[U({\theta}_{a},{\theta}_{b})]$
$\hskip-43.36243pt+\sum_{s\in\{a,b\}}n_{s}(u_{+}\alpha_{s}\bm{\Phi}_{s,(i)}^{1}({\theta_{s}})+u_{+}(1-\alpha_{s})\bm{\Phi}_{s,(i)}^{0}({\theta_{s}})+u_{+}\alpha_{s}\bm{\Psi}_{s,(i)}^{1}({\theta_{s}})-u_{-}(1-\alpha_{s})\bm{\Psi}_{s,(i)}^{0}({\theta_{s}})),$ (7)
@@
[…]
@@
[…]
@@
[…]
The proofs follow from applying similar techniques to those in [Zhang et al., 2022], combined with the results of Lemma 5; these proofs are included in Appendix A.2.3 for completeness.
Lemmas 4-7, together with Proposition 2, fully characterize the possible (Stackelberg) equilibria of the game between agents and the firm, depending on the relative cost-efficacy of strategic actions, whether the firm is cognizant of agents’ strategic behavior, and whether the firm imposes any fairness constraints. We next use these characterizations to assess the impacts of anticipating agents’ strategic behavior on the optimal decision thresholds and the firm’s utility, as well as the fairness implications of this across groups.
@@ 6 Effects of Anticipating Strategic Behavior
We first proceed with investigating the impacts of anticipating agents’ strategic behavior on the optimal policies and the firm’s utility by comparing the non-strategic policy $\theta^{U}_{s}$ (from Lemma 4) with the strategic policy $\hat{\theta}_{s}$ (from Lemma 5) in the absence of any fairness interventions. For concreteness, we do this analysis for cases when all agents opt for the same type of best-response, and specifically, for Types 1 and 3 of agents’ best responses, which can be interpreted as relatively low and high improvement costs, respectively, as illustrated in Section 4. Accordingly, we will also refer to Type 1 and Type 3 equilibria as “affordable improvement” and “costly improvement”, respectively. Notably, Type 3/costly improvement equilibria only include manipulation decisions, whereas Type 1/affordable improvement equilibria include both manipulation and improvement decisions at equilibrium.
Contrasting these two types of equilibria, we can investigate to what extent anticipating gaming can prevent manipulation but also potentially encourage improvement decisions. Notably, while some prior works (e.g., Zhang et al. [2022], Xie et al. [2024b]) show that anticipating gaming can reduce manipulations and potentially encourage improvements, they are limited by the assumption that these actions are only available to unqualified individuals. Our work takes a broader view by considering the availability of these actions to any agent (in addition to other model extensions noted earlier), and therefore sheds light on how these effects differ across qualification states, an aspect not considered in prior work. Our broader view also has implications for the fairness of anticipating strategic behavior, precisely as there are differences in how qualified vs. unqualified agents are impacted by the firm’s anticipation of strategic responses.
@@
Assumption 4 states that the cost and effectiveness of the manipulation action is independent of the qualification state. For instance, using a Generative AI tool to solve basic homework problems has the same cost and effectiveness regardless of a student’s preparation level/qualification state. In contrast, the second part of the assumption implies that the improvement action is more effective for qualified agents. Continuing on the same example, when a qualified student with a stronger background spends time preparing well for an exam, their effort can boost their exam grades more than a less prepared student spending the same time.
Given Assumptions 3 and 4, we can now establish relative orderings between $\theta_{s}$ and $\hat{\theta}_{s}$, and their impact on the different components of the firm’s utility. Recall, from the expression for the firm’s utility (5), that $\bm{\Phi}_{s,(i)}^{y}({\theta_{s}})$ and $\bm{\Psi}_{s,(i)}^{y}({\theta_{s}})$ capture the change in the firm’s utility due to label $y$ agents from group $s$ who opt for improvement and manipulation, respectively, in equilibrium Type (i), given a threshold $\theta_{s}$.
@@
[…]
The proof is provided in Appendix A.3.1. We next provide a brief intuitive interpretation for the findings of this proposition. We discuss the impacts of relaxing the assumptions of Proposition 8 in Appendix A.3.2, as well as in the numerical experiments in Section 7.
The key takeaways behind Proposition 8 is as follows: by anticipating agents’ strategic responses, the firm chooses a higher threshold than a non-strategic firm, consequently reducing incentive toward manipulation for unqualified agents while increasing manipulation incentives for qualified ones, both of which align with improving its utility. Notably, the fact that anticipation of gaming increases the “social burden” on qualified individuals, by forcing them to opt of manipulation has also been observed under different models in prior work [Milli et al., 2019]. Moreover, this higher threshold has even more preferred consequences for the scenario of affordable improvement (Type 1). It increases genuine improvement incentives for all agents, and especially for unqualified agents, for whom genuine improvement becomes the only successful path to acceptance. Together, these observations highlight how anticipating strategic responses will not only reduce gaming of the algorithm by unqualified individuals (as also shown in prior work), but also lead to increased improvement incentives among all agents (a new effect identified by our model). An additional detailed and illustrative interpretation of both costly-improvement (Type 3) and affordable-improvement (Type 1) equilibria, and how each drive the changes in manipulation vs. improvement, is provided in Appendix A.3.3.